1. ( y-2x)dx+xdy=0 ,y(-1)=3
  2. dy=(x-4xy)dx ,y(0)= - 1/4
1
Expert's answer
2021-11-04T19:57:56-0400

Let us solve the following differential equations.


1. "(y-2x)dx+xdy=0 ,y(-1)=3"

This equation can be rewritten in the form "xy'+y-2x=0" or "(xy)'=2x." It follows that "xy=x^2+C," and we conclude that the general solution is of the form: "y=x+\frac{C}x." Then "3=y(-1)=-1-C," and thus "C=-4." The solution of "(y-2x)dx+xdy=0 ,y(-1)=3," is "y=x-\frac{4}x."


2. "dy=(x-4xy)dx ,y(0)= - 1\/4"

This equation is equivalent to "\frac{dy}{1-4y}=xdx," and hence "\int\frac{dy}{1-4y}=\int xdx." Then "-\frac{1}4\ln|1-4y|=\frac{x^2}2+C."

It follows that "y=-\frac{1}4" when "x=0," and hence "-\frac{1}4\ln 2=C."

We conclude that the solution of "dy=(x-4xy)dx ,y(0)= - 1\/4"

is of the form "-\frac{1}4\ln|1-4y|=\frac{x^2}2-\frac{1}4\ln 2."